OpenStudy (anonymous):

Find the slope perpendicular to the line x+7y=-3

OpenStudy (kitkat16):

7y=-x-3 Divide each term in the equation by 7 7y/7= -1/7x -3/7 simplify y=-1/7x-3/7 use y=mx+b m= slope b=y-intercept m= - 1/7 b= - 3/7

OpenStudy (kitkat16):

wait sorry you are look for a\the slope perpendicular

OpenStudy (anonymous):

That's what I got, but I thought you have to use the reciprocal since it's perpendicular so wouldn't the answer be 7?

OpenStudy (kitkat16):

yup its perp so it is 7

OpenStudy (anonymous):

Okay thanks! (:

OpenStudy (kitkat16):

mperp=7 see you had it! Good job!